A relatively quiet early week:
Outstanding issues:Wed May 19 plan:
- Intensities and rates
- Presently in maintenance. Other are Booster RF (early this morning) are 2.5 GeV development on Mon, and the ODH sensor failure early Mon AM.
- Automatic orbit corrections were started Tuesday evening around 6 PM.
- Gap cleaning continues to run since Sat May 15 about 14:23, or fill 13338.
- 2.5 GeV testing results.
- Current end of 7.7 GeV run is scheduled for Thu May 27. Todd departs for IPAC early Thu May 20 (tomorrow).
- The RHIC 2010 web page now has some semi-live performance plots for run tracking.
Beam parameters from Kip
- Maintenance today, APEX tomorrow.
- Keep running physics!
are now continuously on call